### v1

# Unified Performance Control for Non-Square Nonlinear Systems with Relaxed Controllability

Abstract
In this paper, we investigate the problem of unified prescribed performance tracking for a class of non-square strict-feedback nonlinear systems in the presence of actuator faults under relaxed controllability conditions. By using a skillful matrix decomposition and introducing some feasible auxiliary matrices, a more generalized controllability condition than the current state of the art is constructed, which can be applied to both square and non-square nonlinear systems subject to actuator faults and unknown yet time-varying control gain. Incorporating the relaxed controllability conditions and the uniform performance specifications into the backstepping design procedure, a prescribed performance fault-tolerant controller is developed that can achieve different performance demands without modifying the controller structure, which is more flexible and practical. In addition, the destruction of the system stability by unknown auxiliary matrices and unknown nonlinearities is circumvented by embedding the available core information of the state-dependent uncertainties into the design procedure. Both theoretical analysis and numerical simulation demonstrate the effectiveness and benefits of the proposed method.
Index Terms:
Non-square nonlinear systems, controllability condition, prescribed performance, unknown control gain, actuator fault.
I Introduction
Over the past decade, significant progress has been made in the field of adaptive control for uncertain nonlinear systems, with applications spanning robotics [1], autonomous vehicles [2], quadrotors [3], and more. It should be emphasized that the primary challenge prior to control design for systems with uncertain nonlinearities is to establish a suitable controllability condition, that is, to impose appropriate restrictions on the unknown control gains to guide controller design and stability analysis.
For single-input single-output systems, a common controllability condition is to assume, without loss of generality, that there exists a constant lower bound on the control gain $g$. However, such a condition is usually not applicable or conservative for multiple-input multiple-output (MIMO) systems since the control gain is a matrix. In order to solve the control problems of this type of system, some classical assumptions have been imposed in the existing literature. To name a few, one classical assumption imposed on the control gain matrix $\bm{g}$ is requiring that $\bm{g}$ is symmetric and positive definite (SPD) [4], which is mathematically elegant but only applicable to a few practical systems, such as robotic systems [5]. Another improved version of the assumption has been proposed in [6], i.e., assuming that $\bm{g}+\bm{g}^{T}$ is uniformly positive definite, which generally aligns with the physical models of several practical systems, such as robotic systems [5], the wheeled inverted pendulum systems [7], and thus its validity has been widely recognized [8, 9, 10].
However, these controllability conditions (the restrictive assumptions imposed on the control gain matrix) are not always satisfactory. On the one hand, not all practical systems satisfy such conditions, e.g., high-speed train systems [11] and quadrotors [3], since the control gain matrix $\bm{g}$ in their dynamics is neither SPD nor satisfies the fact that $\bm{g}+\bm{g}^{T}$ is SPD. On the other hand, it is known that actuator failure usually destroys this controllability condition. Specifically, when actuators experience multiplicative failures, the original $\bm{g}$ will be right-multiplied by an actuation effectiveness matrix $\bm{\rho}$, as mentioned in [12], the strong controllability condition established in [6] might be invalid. A number of noteworthy efforts have been made to establish more relaxed conditions of controllability. For example, by proposing a sector condition with a design parameter $\bm{K}$, i.e., assuming that $\bm{K}\bm{g}+\bm{g}^{T}\bm{K}^{T}$ is uniformly positive-definite, the control problem of quadrotors can be solved [13]. To account for the destruction of controllability by actuator failures, some works have embedded the actuation effectiveness matrix $\bm{\rho}$ into the controllability condition in [6], i.e., assuming that $\bm{g}\bm{\rho}+\bm{\rho}\bm{g}^{T}$ is uniformly positive definite [14], but it is extremely demanding on the parametric properties of $\bm{g}$ and $\bm{\rho}$, making it hard to satisfy in practice. Alternative approaches have incorporated the actuation effectiveness matrix into the Lyapunov function, requiring the actuation effectiveness matrix to be differentiable with respect to time [12, 15].
Inspired by the above discussion, in this paper, we focus on solving the problem of prescribed performance control (PPC) of strict-feedback MIMO systems under a more relaxed controllability condition. Unlike the existing literatures on PPC for MIMO systems (e.g., [8, 10]), the system considered may have a non-square control gain matrix, and there may be faulty actuators that violate the controllability assumption imposed in these literature. The core technique of this paper is to decompose the control gain matrix and construct several auxiliary (not necessarily known) matrices to reconstruct the controllability of the system in the control design. The considered problem is well addressed by developing a framework for compatibility of this technique with PPC and backstepping. Notice that the proposed approach is different classical PPC since it does not rely on a closed-loop initialisation set and does not require control redesign and reanalysis for different performance requirements. Preliminary result of controllability relaxation for square nonlinear systems is presented in [16], where the classical PPC is considered. In this paper, we further relax the controllability condition for non-square systems subject to actuator faults, while eliminating the restrictions of the classical PPC so that prescribed performance tracking control with a unified framework can be achieved.
The main features and contributions of this work include:
  • The controllability conditions of MIMO systems with actuator faults have been relaxed where the control gain matrices are allowed to be non-square and the control design requires less a priori knowledge about such matrices.
  • By employing practical matrix decomposition and several auxiliary matrices, we propose a design framework applicable to both square and non-square systems, which paves the way for solving the global prescribed performance tracking control problem for MIMO nonlinear systems.
  • It is shown that the proposed controller exhibits better robustness in the sense of actuator failures and uncertain control gains compared to those in the existing literature (see, for instance, [12, 15]). In other words, the developed controller is applicable to the cases with or without actuator failures in the aforementioned literature and vice versa.
[…]
II Problem Statement
Consider a class of high-order MIMO uncertain nonlinear systems in strict-feedback form
$\dot{\bm{x}}_{i}=$ $\bm{f}_{i}(\bm{X}_{i})+\bm{g}_{i}(\bm{X}_{i},t)\bm{x}_{i+1}+{\bm{d}_{i}}(\bm{X}_{i}),i=1,\ldots,N-1,$
$\dot{\bm{x}}_{N}=$ $\bm{f}_{N}(\bm{X}_{N})+\bm{g}_{N}(\bm{X}_{N},t)\bm{u}_{a}+{\bm{d}_{N}}(\bm{X}_{N}),$
$\bm{y}=$ $\bm{x}_{1},$ (1)
where $\bm{x}_{i}=[x_{i1},\ldots,x_{in}]^{T}\in\mathbb{R}^{n}$ ($i=1,\ldots,N$) is the $i$th state vector; $\bm{X}_{i}=[\bm{x}_{1}^{T},\ldots,\bm{x}_{i}^{T}]^{T}\in\mathbb{R}^{n\times i}$; ${\bm{y}}=[y_{1},\ldots,y_{n}]^{T}\in\mathbb{R}^{n}$ is the output vector; ${\bm{f}_{i}}(\bm{X}_{i}):\mathbb{R}^{in}\rightarrow\mathbb{R}^{n}$ is some nonlinear uncertainty; ${\bm{d}_{i}}(\bm{X}_{i}):\mathbb{R}^{in}\rightarrow\mathbb{R}^{n}$ denotes external disturbance; $\bm{u}_{a}=[u_{a1},\ldots,u_{am}]^{T}\in\mathbb{R}^{m}$ is the input vector (the output of actuators); ${\bm{g}_{i}}(\bm{X}_{i}):\mathbb{R}^{in}\rightarrow\mathbb{R}^{n\times m}$, $n\leq m$ is an unknown control gain matrix.
As unanticipated actuator faults may occur for “long-term” operation [17], in such case, the abnormal actuator input-output model can be described as
$\bm{u}_{a}=\bm{\rho}(t)\bm{u}+\bm{\upsilon}(t)$ (2)
where $\bm{u}=[u_{1},\ldots,u_{m}]^{T}\in\mathbb{R}^{m}$ is the designed control input; $\bm{\rho}=\textmd{diag}\{\rho_{1},\ldots,\rho_{m}\}\in\mathbb{R}^{m\times m}$ is the actuation effectiveness matrix that does not need to be piecewise continuous; $\bm{\upsilon}=[\upsilon_{1},\ldots,\upsilon_{m}]^{T}\in\mathbb{R}^{m}$ represents the bounded time varying bias fault, i.e., $\|\bm{\upsilon}\|<\bar{\upsilon}<\infty$ with $\bar{\upsilon}$ being a positive constant. In this subsection, we consider the partial loss of effectiveness (PLOE) case, i.e., $\rho_{j}\in(0,1],j=1,\ldots,m$. Such a scenario may occurs when the attitude control actuators of the aircraft or spacecraft are rotationally limited due to long-term operation [18, 19].
Denote the desired trajectory as $\bm{y}^{*}=[y^{*}_{1},\ldots,y^{*}_{m}]^{T}\in\mathbb{R}^{m}$, and define the tracking error as $\bm{e}=\bm{x}_{1}-\bm{y}^{*}=[e_{1},\ldots,e_{m}]^{T}$. The control objective in this paper is to design a control law such that
  1. 1.
    All signals in the closed loop systems are bounded in the presence of actuator faults;
  2. 2.
    The tracking error $\bm{e}$ is always within the prescribed performance bound, i.e.,
    $\mathcal{H}(-\underline{\delta}_{j}\varphi_{j}(t))<e_{j}(t)<\mathcal{H}(\bar{\delta}_{j}\varphi_{j}(t)),\ j=1,2,\ldots,m,$ (3)
    where $0<\underline{\delta}_{j}$, $\bar{\delta}_{j}\leq 1$ are some positive constants, $\mathcal{H}$ is a time-varying boundary function, and $\varphi_{j}$ denotes a time-varying scaling function. The detailed definitions of $\mathcal{H}$ and $\varphi_{j}$ can be found in Section III-A.
To achieve this objective, the following assumptions are needed.
Assumption 1
The desired tracking trajectory $\bm{y}^{*}$ and its derivatives up to $n$th order are known and bounded. The system state vector $\bm{x}_{i}$ is available for control design.
Assumption 2
The control gain matrix $\bm{g}_{i}$ can be decomposed as
$\bm{g}_{i}=\bm{A}\bm{b}_{i},n<m,\ i=1,\ldots,N$ (4)
where $\bm{A}=[\bm{I}_{n},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{(m-n)}]\in\mathbb{R}^{n\times m}$ is a known constant matrix with $\bm{\Lambda}_{j}=[\iota_{j1},\ldots,\iota_{jn}]^{T}\in\mathbb{R}^{n}$ ($j=1,\ldots,m-n$) being a nonzero vector and $\iota_{jp}\geq 0$ ($p=1,\ldots,n$), and $\bm{{b}}_{i}\in\mathbb{R}^{m\times m}$ is an extension matrix of $\bm{g}_{i}$ which is completely unknown. Without loss of generality, we assume ${\bm{g}_{k}}\in\mathbb{R}^{n\times n}$ is square ($k=1,\ldots,N-1$) and ${\bm{g}_{N}}\in\mathbb{R}^{n\times m}$ with $n<m$ is non-square.
Assumption 3
There exists an unknown positive constant $\bar{g}_{j}$ such that $\|\bm{g}_{i}\|\leq\bar{g}_{i}$ ($i=1,\ldots,N$). Further, for $\bm{g}_{j}$ ($j=1,\ldots,N-1$), there exist an unknown diagonal and positive definite matrix $\bm{P}_{1}(t)\in\mathbb{R}^{n\times n}$ and some unknown symmetric and positive definite matrices $\bm{P}_{k}(\bm{X}_{k-1},t)\in\mathbb{R}^{n\times n}$ ($k=2,\ldots,N-1$) such that $\bm{G}_{j}:={\bm{P}_{j}\bm{g}_{j}+\bm{g}_{j}^{T}\bm{P}_{j}}$ ($j=1,\ldots,N-1$) are uniformly sign-definite with known signs, i.e., $0<\lambda_{i}\leq{\underline{\sigma}}(\bm{G}_{i}/2)$, where $\lambda_{i}$ is an unknown constant; and for $\bm{g}_{N}$, there exists an unknown symmetric and positive definite matrix $\bm{P}_{N}(\bm{X}_{N-1},t)\in\mathbb{R}^{n\times n}$ such that $\bm{G}_{N}:={\bm{P}_{N}\bm{{g}}_{N}\bm{\rho}\bm{A}^{T}+\bm{A}\bm{\rho}\bm{{g}}_{N}^{T}\bm{P}_{N}}$ is uniformly sign-definite with known sign, i.e., $0<\lambda_{N}\leq{\underline{\sigma}}(\bm{G}_{N}/2)/\|\bm{A}\|$, where $\lambda_{N}$ is an unknown constant. Without loss of generality, we assume that all $\bm{G}_{i}$, $i=1,\ldots,N$, are uniformly positive definite.
Remark 1
Assumption 3 ensures the controllability of system (1), which is milder than the current state-of-the-art [6, 13, 10, 20, 16]. In [13] and [20], similar conditions are considered, yet $\bm{P}_{i}$ is required to be known and diagonal form, respectively, and the condition in [16] is only available for square systems. Thus, the controllability conditions in [6, 13, 16] can be viewed as some special cases of Assumption 3 since $\bm{P}_{i}$ is allowed to be unknown and can handle non-square nonlinear systems in our case. Moreover, it is worth noting that the auxiliary matrix $\bm{P}_{i}$ is only required for stability analysis, as opposed to needing to be present in the control design as in [13], which thus benefits the implementation of our method.
Remark 2
In contrast to the decomposition of $\bm{g}_{N}$ in [2] and [9], no information about $\bm{g}_{N}$ except the structural features is required in Assumption 2, and the decomposed known matrix $\bm{A}$ is skilfully constructed to be uniform and simple for all systems, and does not prevent the attainment of the controllability condition. In [2] and [9], the controllability condition is established by decomposing the control gain matrix to obtain a system information-based matrix $\bm{L}\in\mathbb{R}^{n\times m}$ with full-row rank, however, such a decomposition is complex and hard to implement for some systems, moreover, such a controllability condition may be invalidated due to the presence of actuator faults (see Example 2 for detail). It is worth emphasizing that by Assumptions 2 and 3, the controllability of the system is ensured by matching the auxiliary matrix $\bm{P}_{i}$ analytically, rather than by solving the matrix $\bm{L}$ in a complex offline manner, as seen in our later development.
In what follows, we present two examples to shed light on the merits of Assumption 2 and Assumption 3.
Example 1: (square case) Consider
$\bm{g}=\begin{bmatrix}2&1\\ 5&3\end{bmatrix},\ \textrm{and}\ \bm{g}+\bm{g}^{T}=\begin{bmatrix}4&6\\ 6&6\end{bmatrix}.$
It is trivial to verify that neither $\bm{g}$ nor $\bm{g}+\bm{g}^{T}$ are SPD. One can choose an auxiliary matrix $\bm{P}$ as follows
$\bm{P}=\begin{bmatrix}3&0\\ 0&1\end{bmatrix},\ \textrm{then}\ \bm{P}\bm{g}+\bm{g}^{T}\bm{P}=\begin{bmatrix}12&8\\ 8&6\end{bmatrix},$
which shows that Assumption 3 is satisfied and the controllability of the system is guaranteed.
Example 2: (non-square case) Consider a spacecraft equipped with four reaction wheels (RWs) [19], and the control gain matrix is defined as $\bm{g}=\bm{J}^{-1}\bm{D}\in\mathbb{R}^{3\times 4}$, where the inertia matrix $\bm{J}\in\mathbb{R}^{3\times 3}$ and configuration matrix $\bm{D}\in\mathbb{R}^{3\times 4}$ are given by
$\bm{J}=\begin{bmatrix}20&1.2&0.9\\ 1.2&5&1.4\\ 0.9&1.4&5\end{bmatrix}\textrm{kg}\cdot\textrm{m}^{2},\ \bm{D}=\begin{bmatrix}1&0&0&{\sqrt{3}}\\ 0&1&0&{\sqrt{3}}\\ 0&0&1&{\sqrt{3}}\end{bmatrix}.$
Assume that the actuation effectiveness matrix is set as $\bm{\rho}(t)=\textmd{diag}\{0.01,0.1+0.01\sin(t),0.01,0.1+0.01\cos(t)\}$. If we choose $\bm{L}=\bm{D}$ as in [21] and [19], by a simple calculation, it can be verified that $\bm{G}^{*}={\bm{{g}}\bm{\rho}\bm{D}^{T}+\bm{D}\bm{\rho}\bm{{g}}^{T}}$ is not positive definite, but we can still make Assumption 3 hold by choosing
$\bm{A}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&1\end{bmatrix}\ \textrm{and}\ \bm{P}=\begin{bmatrix}0.7+0.1\sin(t)&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}.$
Assumption 4
There exist some unknown positive constants $a_{fi}$, $a_{11}$, $a_{12}$, $a_{k1}$, $a_{k2}$ and some known nonnegative scalar “core functions” $\phi_{fi}(\bm{X}_{i})$, $\phi_{k1}(\bm{X}_{k-1})$ and $\phi_{k2}(\bm{X}_{k})$ ($i=1,2,\ldots,N$, $k=2,\ldots,N$) such that
$\|\bm{f}_{i}(\bm{X}_{i})+{\bm{d}_{i}}(\bm{X}_{i})\|\leq a_{fi}\phi_{fi}(\bm{X}_{i}),$ (5)
$\|\bm{P}_{1}(t)\|\leq a_{11},\ \big{\|}\frac{\partial\bm{P}_{1}(t)}{\partial t}\big{\|}\leq a_{12},$ (6)
$\|\bm{P}_{k}(\bm{X}_{k-1},t)\|\leq a_{k1}\phi_{k1}(\bm{X}_{k-1}),$
$\big{\|}\frac{\partial\bm{P}_{k}(\bm{X}_{k-1},t)}{\partial t}\big{\|}\leq a_{k2}\phi_{k2}(\bm{X}_{k}),$ (7)
where $\phi_{fi}$, $\phi_{k1}$ and $\phi_{k2}$ are radially unbounded.
Remark 3
Assumption 4 is not a stringent requirement, and many practical systems can readily fulfill such a condition by extracting crude model information [9, 15, 2]. Particularly, even if no ”core function” can be extracted, one can judiciously choose $\phi_{fi}=\phi_{k1}=\phi_{k2}=1$, which is equivalent to impose some unknown upper bounds on $\|\bm{f}_{i}\|$, $\|\bm{d}_{i}\|$, $\|\bm{P}_{k}\|$ and $\|{\partial\bm{P}_{k}}/{\partial t}\|$. Consequently, by employing core functions, non-parametric uncertainties can be handled elegantly with less restrictions.
III Main Result
III-A Non-monotonic Performance Function
To achieve the prescribed tracking performance, we first introduce the rate function $\beta(t)$ satisfying the following conditions: 1) $\beta(0)=1$; 2) $\beta(t)\in[0,1),\forall t>0$; and 3) $\beta^{(i)},i=0,1,...,N$, is known, bounded, and piecewise continuous.
Based on the rate function, we construct the following scaling function
$\varphi(t)=(\varphi_{0}-\varphi_{f})\beta(t)+\varphi_{f},$ (8)
where $\varphi_{f}$ and $\varphi_{0}$ are some design parameters and satisfy $0<\varphi_{f}<\varphi_{0}\leq 1$. It is readily shown that $\varphi(t)$ exhibits the following properties: 1) $\varphi(0)=\varphi_{0}$ and $\varphi:(0,\infty)\rightarrow[\varphi_{f},\varphi_{0})$ for $t>0$; and 2) $\varphi^{(i)},i=0,1,...,n,$ is bounded, known, and piecewise continuous over the interval $[0,\infty)$.
Definition 1: A continuous function $\mathcal{H}:(-b_{0},b_{0})\rightarrow[-\infty,\infty)$ is called a intermediate function if 1) $\mathcal{H}$ is continuously differentiable; 2) $\mathcal{H}(\varsigma)=\mathcal{H}(-\varsigma)$, $\mathcal{H}(0)=0$, and $\lim_{\varsigma\rightarrow\pm b_{0}}\mathcal{H}(\varsigma)=\pm\infty$ ; and 3) $0<\upsilon\leq\dot{\mathcal{H}}$ and $\lim_{\varsigma\rightarrow\pm b_{0}}\dot{\mathcal{H}}(\varsigma)=\infty$, where $\upsilon$ is a constant.
It is worth noting that there are many examples of such $\mathcal{H}$, e.g., $\mathcal{H}(\varsigma)=\frac{\varsigma}{\sqrt{1-\varsigma^{2}}}$ and $\mathcal{H}(\varsigma)=\tan(\frac{\pi}{2}\varsigma)$ with $b_{0}=1$.
According to the definitions of $\varphi$ and $\mathcal{H}$, we call $\mathcal{H}(\varphi)$ the performance function, and it is trivial to prove that $\mathcal{H}$ is a non-monotonic function as $\varphi$ is a non-monotonic function. To facilitate the presentation, this paper adopts the non-monotonic performance function as in [15]
$\mathcal{H}(\varphi)=\frac{l\varphi}{\sqrt{1-\varphi^{2}}}\ \textrm{with}\ \beta(t)=\exp(-\gamma t)\cos^{2}(t),$ (9)
where $l$ and $\gamma$ are some positive constants. Note that in [8, 12], the scaling function $\varphi(t)$ is monotonically decreasing in time, however, as mentioned in [22] and [15], utilizing a non-monotonic function that can widen the performance boundary over certain time intervals may be beneficial in some practical scenarios.
III-B Error Transformation
To achieve the performance constraint (3)) for $t\geq 0$, we define the error transformation function as
${s}_{j}(t)=\frac{\zeta_{j}}{(\underline{\delta}_{j}+\zeta_{j})(\bar{\delta}_{j}-\zeta_{j})},j=1,2,\ldots,n$ (10)
with $\zeta_{j}=\frac{\eta_{j}}{\varphi_{j}}$ and $\eta_{j}=\frac{e_{j}}{\sqrt{e_{j}^{2}+l_{j}^{2}}}$.
It is seen from (10) that $s_{j}$ has the following properties: 1) for any $-\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$, $s_{j}\rightarrow\pm\infty$ as $\zeta_{j}\rightarrow\bar{\delta}_{j}$ or $\zeta_{j}\rightarrow-\underline{\delta}_{j}$; 2) if $\zeta_{j}(0)$ satisfies $-\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$ and $s_{j}(t)$ is bounded $\forall t\geq 0$, then there exist some constants $-\underline{\delta}_{1j}$ and $\bar{\delta}_{1j}$ so that $-\underline{\delta}_{j}<-\underline{\delta}_{1j}\leq\zeta_{j}(0)\leq\bar{\delta}_{1j}<\bar{\delta}_{j}$.
The following lemma shall be used to explain how (3) is ensured from (10).
Lemma 1
[15] If the control to be developed is able to guarantee the boundedness of $s_{j}(t)$ for $t\geq 0$, then for any $\zeta_{j}(0)$ satisfying $-\underline{\delta}_{j}<\zeta_{j}(0)<\bar{\delta}_{j}$, (3) is ensured.
Proof: The detailed proof is provided in [15], thus omitted.
The notable advantage of this error transformation is that the unified system framework for prescribed asymmetric and symmetric performance can be achieved by choosing the design parameters $\underline{\delta}_{j}$, $\bar{\delta}_{j}$ and $\varphi(0)$ (see [15] for details). Specifically, this error transformation framework can achieve global results similar to [22, 23, 24] (without considering error overshoot), as well as results similar to PPC [8, 20] and asymmetric performance behaviors.
III-C System Transformation
Denote ${\bm{\varepsilon}}_{1}=[{s}_{1},\cdots,{s}_{n}]^{T}$ as the transformed tracking error where ${s}_{j}$ ($j=1,\ldots,n$) is defined in (10). Differentiating (10) w.r.t. time, we have
$\dot{s}_{j}=\mu_{j}\dot{\zeta}_{j}=\frac{\mu_{j}r_{j}}{\varphi_{j}}\dot{e}_{j}-\frac{\mu_{j}\dot{\varphi}_{j}\eta_{j}}{\varphi_{j}^{2}}=w_{j}(\dot{e}_{j}+v_{j})$ (11)
where $\mu_{j}=\frac{\underline{\delta}_{j}\bar{\delta}_{j}+\zeta_{j}^{2}}{(\underline{\delta}_{j}+\zeta_{j})^{2}(\bar{\delta}_{j}-\zeta_{j})^{2}}$ is well-defined in the set $\Omega_{\zeta_{j}}=\{\zeta_{j}\in\mathbb{R}:-\underline{\delta}_{j}<\zeta_{j}<\bar{\delta}_{j}\}$, $r_{j}=\frac{l_{j}^{2}}{\sqrt{e_{j}^{2}+l_{j}^{2}}(e_{j}^{2}+l_{j}^{2})}$, $w_{j}=\frac{\mu_{j}r_{j}}{\varphi_{j}}$, and $v_{j}=-\frac{\dot{\varphi}_{j}\eta_{j}}{\varphi_{j}r_{j}}$. Further, (11) can be written in compact form as
$\dot{\bm{\varepsilon}}_{1}=\bm{W}(\dot{\bm{e}}+\bm{V})$ (12)
where $\bm{W}=\textrm{diag}\{w_{j}\}\in\mathbb{R}^{n\times n}$ and $\bm{V}=[v_{1},\ldots,v_{n}]^{T}\in\mathbb{R}^{n}$. Then, from the system (1), we obtain
$\dot{\bm{\varepsilon}}_{1}=\bm{W}\left(\bm{f}_{1}({\bm{x}}_{1})+\bm{g}_{1}({\bm{x}}_{1},t)\bm{x}_{2}+{\bm{d}_{1}}({\bm{x}_{1}},t)-\dot{\bm{y}}_{d}+\bm{V}\right),$ (13)
where $\dot{\bm{\varepsilon}}_{1}=[\dot{s}_{1},\cdots,\dot{s}_{n}]^{T}$. According to the properties of the transformations (10), $\bm{W}$ and $\bm{V}$ are available for control design, moreover, $\bm{W}$ is invertible and $\bm{W}>0$ as long as $-\underline{\delta}_{j}<\zeta_{j}<\bar{\delta}_{j}$ over the interval $t\in[0,\tau_{\textrm{max}}]$. Now the original system (1) can be transformed as
$\dot{\bm{\varepsilon}}_{1}$ $=\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}\bm{x}_{2}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right)$
$\dot{\bm{x}}_{i}$ $=\bm{f}_{i}+\bm{g}_{i}\bm{x}_{i+1}+{\bm{d}_{i}},i=2,\ldots,N-1,$
$\dot{\bm{x}}_{N}$ $=\bm{f}_{N}+\bm{g}_{N}(\bm{\rho}\bm{u}+\bm{\upsilon})+{\bm{d}_{N}}.$ (14)
Therefore, (14) is still a strict-feedback system since $\bm{W}>0$, and hence, we can achieve the prescribed performance control of the original system (1) by designing the controller to stabilize the transformed system (14).
III-D Controller Design
To stabilize system (14), here we develop a robust adaptive controller by applying the backstepping method in [25]. Let us begin the control design by defining the virtual control errors as
${\bm{\varepsilon}}_{i}=\bm{x}_{i}-{\bm{a}}_{i-1},\ i=2,\ldots,N.$ (15)
where ${\bm{a}}_{i-1}$ is a virtual controller to be designed at the $i-1$th step and the actual controller $\bm{u}$ is derived at the final step.
Step 1: From (14) and (15), it is derived that
$\dot{\bm{\varepsilon}}_{1}=\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}{\bm{\varepsilon}}_{2}+\bm{g}_{1}{\bm{a}}_{1}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right),$ (16)
Then, consider the Lyapunov candidate function as
$V_{11}=\frac{1}{2}{\bm{\varepsilon}}_{1}^{T}\bm{P}_{1}{\bm{\varepsilon}}_{1},$ (17)
where $\bm{P}_{1}\in\mathbb{R}^{n\times n}$ is an unknown diagonal and positive definite matrix under Assumption 3. Taking the time derivative of $V_{11}$ along (16) yields
$\dot{V}_{11}=$ ${\bm{\varepsilon}}_{1}^{T}\bm{P}_{1}\bm{W}\left(\bm{f}_{1}+\bm{g}_{1}{\bm{\varepsilon}}_{2}+\bm{g}_{1}{\bm{a}}_{1}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right)$ (18)
$+\frac{1}{2}{\bm{\varepsilon}}_{1}^{T}\dot{\bm{P}}_{1}{\bm{\varepsilon}}_{1}$
$=$ ${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{a}}_{1}+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}+\bm{h}_{1}$
with $\bm{h}_{1}={\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\left({\bm{f}}_{1}+{\bm{d}_{1}}-\dot{\bm{y}}_{d}+\bm{V}\right)+\frac{1}{2}{\bm{\varepsilon}}_{1}^{T}\dot{\bm{P}}_{1}{\bm{\varepsilon}}_{1}$ where the diagonal property of $\bm{W}$ and $\bm{P}_{1}$ is used. Upon employing Young’s inequality along with Assumption 3 and 4, one has
${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}({\bm{f}}_{1}+{\bm{d}_{1}})\leq$ $\ \|\bm{W}\bm{\varepsilon}_{1}\|^{2}a_{11}^{2}a_{f1}^{2}\phi_{11}^{2}\phi_{f1}^{2}+\frac{1}{4},$
${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}(-\dot{\bm{y}}_{d}+\bm{V})\leq$ $\ \|\bm{W}\bm{\varepsilon}_{1}\|^{2}a_{11}^{2}\phi_{11}^{2}(\|\dot{\bm{y}}_{d}\|+\|\bm{V}\|)^{2}$
$+\frac{1}{4},$
Since $\frac{1}{2}{\bm{\varepsilon}}_{1}^{T}\dot{\bm{P}}_{1}{\bm{\varepsilon}}_{1}\leq\frac{1}{2}\|\bm{W}\bm{\varepsilon}_{1}\|^{2}a_{12}(\lambda_{\textrm{min}}(\bm{W}))^{-2}\phi_{12}$, then the uncertain function $\bm{h}_{1}$ can be upper bounded by
$\bm{h}_{1}\leq\|\bm{W}\bm{\varepsilon}_{1}\|^{2}\theta_{1}\Phi_{1}+\frac{1}{2}$ (19)
with $\theta_{1}=\max\{a_{11}^{2}a_{f1}^{2},a_{11}^{2},a_{12}\}$ being an unknown positive constant and $\Phi_{1}=\phi_{11}^{2}\phi_{f1}^{2}+\phi_{11}^{2}(\|\dot{\bm{y}}_{d}\|+\|\bm{V}\|)^{2}+\frac{1}{2}(\lambda_{\textrm{min}}(\bm{W}))^{-2}\phi_{12}$ denoting a known and computable scalar function. Then, we have
$\dot{V}_{11}\leq$ $\ {\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}({\bm{a}}_{1}+{\bm{\varepsilon}}_{2})+\|\bm{W}\bm{\varepsilon}_{1}\|^{2}\theta_{1}\Phi_{1}+\frac{1}{2}.$ (20)
At this stage, we construct the virtual controller $\bm{a}_{1}$ as:
$\bm{a}_{1}$ $=-\kappa_{1}\bm{W}{\bm{\varepsilon}}_{1}-{\hat{\theta}_{1}{\Phi_{1}}{\bm{W}{\bm{\varepsilon}}_{1}}},$ (21)
$\dot{\hat{\theta}}_{1}$ $={\sigma_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}}\Phi_{1}-\mu_{1}\hat{\theta}_{1},\ \hat{\theta}_{1}(0)\geq 0,$ (22)
where $\kappa_{1}>0$, $\mu_{1}>0$ and $\sigma_{1}>0$ are design parameters, $\hat{\theta}_{1}$ is the estimate of unknown constant $\theta_{1}$. Now we consider the complete Lyapunov function candidate as
$V_{1}=V_{11}+\frac{1}{2\lambda_{1}\sigma_{1}}{\tilde{\theta}_{1}}^{2}.$ (23)
where $\tilde{\theta}_{1}=\theta_{1}-\lambda_{1}\hat{\theta}_{1}$ denotes the parameter estimate error. Differentiating (23) and using (20)-(22) yields
$\dot{V}_{1}\leq$ $-{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}\left(\kappa_{1}\bm{W}{\bm{\varepsilon}}_{1}+{\hat{\theta}_{1}{\Phi_{1}}{\bm{W}{\bm{\varepsilon}}_{1}}}\right)+\frac{\mu_{1}}{\sigma_{1}}\tilde{\theta}_{1}\hat{\theta_{1}}$ (24)
$+(\theta_{1}-\tilde{\theta}_{1}){\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}\Phi_{1}}+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}+\frac{1}{2}.$
Note that
$\bm{P}_{1}\bm{g}_{1}=\,\frac{\bm{G}_{1}}{2}+\frac{\left({\bm{P}_{1}\bm{g}_{1}-\bm{g}_{1}^{T}\bm{P}_{1}}\right)}{2},$ (25)
where $\bm{G}_{1}$ is a symmetric matrix and ${\bm{P}_{1}\bm{g}_{1}-\bm{g}_{1}^{T}\bm{P}_{1}}$ is a skew-symmetric matrix, then by Assumption 3, it holds that
${\bm{\varepsilon}}_{1}^{T}\bm{W}\left({\bm{P}_{1}\bm{g}_{1}-\bm{g}_{1}^{T}\bm{P}_{1}}\right)\bm{W}{\bm{\varepsilon}}_{1}=0,$ (26)
${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{G}_{1}\bm{W}{\bm{\varepsilon}}_{1}\geq\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}.$
Note that in (LABEL:sgs12), the potential destruction of the controllability condition by the performance incidental matrix $\bm{W}$ is neatly avoided by extending the ${\bm{\varepsilon}}_{1}$-based feedback to ($\bm{W}{\bm{\varepsilon}}_{1}$)-based feedback. Subsequently, inserting (25) and (LABEL:sgs12) into (24) yields
$\dot{V}_{1}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\lambda_{1}\hat{\theta}_{1}{\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}\Phi_{1}}+\frac{\mu_{1}}{\sigma_{1}}\tilde{\theta}_{1}\hat{\theta_{1}}$ (27)
$+(\theta_{1}-\tilde{\theta}_{1}){\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}\Phi_{1}}+\frac{1}{2}+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}$
$\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}+\Delta_{1}+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2},$
where the fact that $\frac{\mu_{1}}{\sigma_{1}}\tilde{\theta}_{1}\hat{\theta}_{1}\leq\frac{\mu_{1}}{2\sigma_{1}}{\theta}_{1}^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}$ is used, and $\Delta_{1}=\frac{\mu_{1}}{2\sigma_{1}}{\theta}_{1}^{2}+\frac{1}{2}$ is a positive constant. The term ${\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}$ will be handled in next step.
Step 2: The time derivative of ${\bm{\varepsilon}}_{2}=\bm{x}_{2}-\bm{a}_{1}$ is
$\dot{\bm{\varepsilon}}_{2}=\bm{f}_{2}+\bm{g}_{2}{\bm{\varepsilon}}_{3}+\bm{g}_{2}{\bm{a}}_{2}+{\bm{d}_{2}}-\dot{\bm{a}}_{1},$ (28)
where $\dot{\bm{a}}_{1}=\frac{\partial\bm{a}_{1}}{\partial\bm{x}_{1}}\left(\bm{f}_{1}+\bm{g}_{1}\bm{x}_{2}+{\bm{d}_{1}}\right)+\bm{\omega}_{1}$ with $\bm{\omega}_{1}=\sum_{k=0}^{1}\big{(}\frac{\partial\bm{a}_{1}}{\partial\bm{y}_{d}^{(k)}}{\bm{y}}_{d}^{(k+1)}+\frac{\partial\bm{a}_{1}}{\partial{\varphi}^{(k)}}{\varphi}^{(k+1)}\big{)}+\frac{\partial\bm{a}_{1}}{\partial{\hat{\theta}}_{1}}\dot{\hat{\theta}}_{1}$.
Choose the Lyapunov candidate function as $V_{21}=V_{1}+\frac{1}{2}{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}{\bm{\varepsilon}}_{2}$, where $\bm{P}_{2}\in\mathbb{R}^{n\times n}$ is defined in Assumption 3. Taking the time derivative of $V_{21}$ along (28) yields
$\dot{V}_{21}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}+\Delta_{1}$ (29)
$+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{a}}_{2}+\bm{h}_{2}$
where $\bm{h}_{2}={\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\left(\bm{f}_{2}+{\bm{d}_{2}}-\frac{\partial\bm{a}_{1}}{\partial\bm{x}_{1}}\left(\bm{f}_{1}+{\bm{d}}_{1}+\bm{g}_{1}\bm{x}_{2}\right)-\bm{\omega}_{1}\right)+{\bm{\varepsilon}}_{1}^{T}\bm{W}\bm{P}_{1}\bm{g}_{1}{\bm{\varepsilon}}_{2}+\frac{1}{2}{\bm{\varepsilon}}_{2}^{T}\dot{\bm{P}_{2}}{\bm{\varepsilon}}_{2}$. Similar to (19), by using Young’s inequality and Assumption 4, it is readily shown that the uncertain function $\bm{h}_{2}$ can be upper bounded by
$\bm{h}_{2}\leq\|\bm{\varepsilon}_{2}\|^{2}\theta_{2}\Phi_{2}+\frac{5}{4}$ (30)
with $\theta_{2}=\max\{a_{21}^{2}a_{f2}^{2},a_{21}^{2}a_{f1}^{2},a_{21}^{2}\bar{g}_{1}^{2},a_{21}^{2},a_{11}^{2}\bar{g}_{1}^{2},a_{22}\}$ being an unknown positive constant and $\Phi_{2}=\phi_{21}^{2}\phi_{f2}^{2}+\phi_{21}^{2}\phi_{f1}^{2}\|\frac{\partial\bm{a}_{1}}{\partial\bm{x}_{1}}\|^{2}+\phi_{21}^{2}\|\frac{\partial\bm{a}_{1}}{\partial\bm{x}_{1}}\|^{2}+\phi_{21}^{2}\|\bm{\omega}_{1}\|^{2}+\phi_{11}^{2}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}+\frac{1}{2}\phi_{22}$ denoting a known and computable scalar function. Then substituting (30) into (29), it follows that
$\dot{V}_{21}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}$ (31)
$+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{a}}_{2}+\|\bm{\varepsilon}_{2}\|^{2}\theta_{2}\Phi_{2}+\Delta_{1}+\frac{5}{4}.$
Now the virtual controller $\bm{a}_{2}$ is designed as:
$\bm{a}_{2}$ $=-\kappa_{2}{\bm{\varepsilon}}_{2}-{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}},$ (32)
$\dot{\hat{\theta}}_{2}$ $={\sigma_{2}\|{\bm{\varepsilon}}_{2}\|^{2}}\Phi_{2}-\mu_{2}\hat{\theta}_{2},\ \hat{\theta}_{2}(0)\geq 0,$ (33)
where $\kappa_{2}>0$, $\mu_{2}>0$ and $\sigma_{2}>0$ are design parameters, $\hat{\theta}_{2}$ is the estimate of $\theta_{2}$.
Constructing the complete Lyapunov function candidate as $V_{2}=V_{21}+\frac{1}{2\lambda_{2}\sigma_{2}}{\tilde{\theta}_{2}}^{2}$, where $\tilde{\theta}_{2}=\theta_{2}-\lambda_{2}\hat{\theta}_{2}$ is the estimate error. Then, differentiating $V_{2}$ and using (31)-(33) yields
$\dot{V}_{2}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\frac{\mu_{1}}{2\sigma_{1}}\tilde{\theta}_{1}^{2}+\lambda_{2}\hat{\theta}_{2}{{\|{\bm{\varepsilon}}_{2}\|^{2}\Phi_{2}}}$ (34)
$+\frac{\mu_{2}}{\sigma_{2}}\tilde{\theta}_{2}\hat{\theta}_{2}-{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}\left(\kappa_{2}{\bm{\varepsilon}}_{2}+{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}}\right)$
$+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}+\Delta_{1}+\frac{5}{4}.$
Similar to the analysis in (25) and (LABEL:sgs12), we have that
${\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}(\kappa_{2}{\bm{\varepsilon}}_{2}+{\hat{\theta}_{2}{\Phi_{2}}{{\bm{\varepsilon}}_{2}}})\geq(\kappa_{2}\lambda_{2}+\lambda_{2}{\hat{\theta}_{2}{\Phi_{2}}})\|{\bm{\varepsilon}}_{2}\|^{2},$ (35)
which further leads to
$\dot{V}_{2}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\kappa_{2}\lambda_{2}\|{\bm{\varepsilon}}_{2}\|^{2}-\sum_{k=1}^{2}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}$ (36)
$+\Delta_{2}+{\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}.$
where the fact that $\frac{\mu_{2}}{\sigma_{2}}\tilde{\theta}_{2}\hat{\theta}_{2}\leq\frac{\mu_{2}}{2\sigma_{2}}{\theta}_{2}^{2}-\frac{\mu_{2}}{2\sigma_{2}}\tilde{\theta}_{2}^{2}$ is used, and $\Delta_{2}=\Delta_{1}+\frac{\mu_{2}}{2\sigma_{2}}{\theta}_{2}^{2}+\frac{5}{4}$ is a positive constant. The term ${\bm{\varepsilon}}_{2}^{T}\bm{P}_{2}\bm{g}_{2}{\bm{\varepsilon}}_{3}$ will be handled in next step.
Step i ($i=3,\ldots,N-1$): Based on the first two steps and considering the Lyapunov function candidate as
$V_{i}=V_{i-1}+\frac{1}{2}{\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}{\bm{\varepsilon}}_{i}+\frac{1}{2\lambda_{i}\sigma_{i}}{\tilde{\theta}_{i}}^{2},\ i=3,\ldots,N-1,$ (37)
where $\bm{P}_{i}\in\mathbb{R}^{n\times n}$ is defined in Assumption 3, and ${\tilde{\theta}_{i}}=\theta_{i}-\lambda_{i}\hat{\theta}_{i}$. It is worth noting that the auxiliary matrix $\bm{P}_{i}$ is $(\bm{X}_{i-1},t)$-dependent rather than $(\bm{X}_{i},t)$-dependent, because $\bm{P}_{i}$ being $(\bm{X}_{i},t)$-dependent would cause $\partial\bm{P}_{i}/\partial t$ to be $(\bm{X}_{i+1},t)$-dependent, and thus the virtual controller $\bm{\alpha}_{i}$ will be $(\bm{X}_{i+1},t)$-dependent, which will lead to the algebraic loop problem in the recursive design procedure.
The virtual controllers $\bm{a}_{i}$ ($i=3,\ldots,N-1$) and adaptive laws $\dot{\hat{\theta}}_{i}$ ($i=3,\ldots,N-1$) can be recursively obtained by following the standard backstepping procedure, as summarized in Table I. Note that $\kappa_{i}>0$, $\mu_{i}>0$ and $\sigma_{i}>0$ are design parameters, $\hat{\theta}_{i}$ is the estimate of $\theta_{i}$ with $\theta_{i}=\max\{a_{i1}^{2}a_{fi}^{2},a_{i1}^{2}a_{f(i-1)}^{2},a_{i1}^{2}\bar{g}_{i-1}^{2},a_{i1}^{2},a_{(i-1)1}^{2}\bar{g}_{i-1}^{2},a_{i2}\}$.
TABLE I: Adaptive backstepping virtual controller.
  Virtual Control Schemes: ($i=3,\ldots,N-1$)
$\ \bm{a}_{i}=-\kappa_{i}{\bm{\varepsilon}}_{i}-{\hat{\theta}_{i}{\Phi_{i}}{{\bm{\varepsilon}}_{i}}},$ (38)
$\ \dot{\hat{\theta}}_{i}={\sigma_{i}\|{\bm{\varepsilon}}_{i}\|^{2}}\Phi_{i}-\mu_{i}\hat{\theta}_{i},\ \hat{\theta}_{i}(0)\geq 0,$ (39)
with $\Phi_{i}=\phi_{i1}^{2}\phi_{fi}^{2}+\phi_{i1}^{2}\phi_{f(i-1)}^{2}\|\frac{\partial\bm{a}_{i-1}}{\partial\bm{x}_{i-1}}\|^{2}+\phi_{i1}^{2}\|\frac{\partial\bm{a}_{i-1}}{\partial\bm{x}_{i-1}}\|^{2}+\phi_{i1}^{2}\|\bm{\omega}_{i-1}\|^{2}+\phi_{(i-1)1}^{2}\|{\bm{\varepsilon}}_{i-1}\|^{2}+\frac{1}{2}\phi_{i2}$,
$\bm{\omega}_{i-1}=\sum_{k=0}^{i-1}\big{(}\frac{\partial\bm{a}_{i-1}}{\partial\bm{y}_{d}^{(k)}}{\bm{y}}_{d}^{(k+1)}+\frac{\partial\bm{a}_{i-1}}{\partial{\varphi}^{(k)}}{\varphi}^{(k+1)}\big{)}+\sum_{k=1}^{i-1}\frac{\partial\bm{a}_{i-1}}{\partial{\hat{\theta}}_{k}}\dot{\hat{\theta}}_{k}$.
Similar to the analysis in Step 2, it is directly deduced from Table I that
$\dot{V}_{i}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\sum_{k=2}^{i}\kappa_{k}\lambda_{k}\|{\bm{\varepsilon}}_{k}\|^{2}$ (40)
$-\sum_{k=1}^{i}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}+\Delta_{i}+{\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}\bm{g}_{i}{\bm{\varepsilon}}_{i+1},$
where $\Delta_{i}=\Delta_{i-1}+\frac{\mu_{i}}{2\sigma_{i}}{\theta}_{i}^{2}+\frac{5}{4}$ is a positive constant, and ${\bm{\varepsilon}}_{i}^{T}\bm{P}_{i}\bm{g}_{i}{\bm{\varepsilon}}_{i+1}$ will be handled in final step.
Step N : From (14) and (15), the time derivative of $\bm{\varepsilon}_{N}$ can be obtained as
$\dot{\bm{\varepsilon}}_{N}=\bm{f}_{N}+\bm{g}_{N}\bm{\rho}\bm{u}+\bm{g}_{N}\bm{v}+{\bm{d}_{N}}-\dot{\bm{a}}_{N-1},$ (41)
where $\dot{\bm{a}}_{N-1}=\sum_{k=1}^{N-1}\frac{\partial\bm{a}_{N-1}}{\partial\bm{x}_{k}}\left(\bm{f}_{k}+{\bm{d}_{k}}+\bm{g}_{k}\bm{x}_{k+1}\right)+\bm{\omega}_{N-1}$ with $\bm{\omega}_{N-1}=\sum_{k=0}^{N-1}\big{(}\frac{\partial\bm{a}_{N-1}}{\partial\bm{y}_{d}^{(k)}}{\bm{y}}_{d}^{(k+1)}+\frac{\partial\bm{a}_{N-1}}{\partial{\varphi}^{(k)}}{\varphi}^{(k+1)}\big{)}+\sum_{k=1}^{N-1}\frac{\partial\bm{a}_{N-2}}{\partial{\hat{\theta}}_{k}}\dot{\hat{\theta}}_{k}$.
Then considering the Lyapunov function candidate as
$V_{N1}=V_{N-1}+\frac{1}{2}{\bm{\varepsilon}}_{N}^{T}\bm{P}_{N}{\bm{\varepsilon}}_{N},$ (42)
where $\bm{P}_{N}\in\mathbb{R}^{n\times n}$ is defined in Assumption 3.
Taking the time derivative of $V_{N1}$ along (41) , and from Assumption 3, it holds that
$\dot{V}_{N1}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\sum_{k=2}^{N-1}\kappa_{k}\lambda_{k}\|{\bm{\varepsilon}}_{k}\|^{2}-\sum_{k=1}^{N-1}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}$ (43)
$+\Delta_{N-1}+{\bm{\varepsilon}}_{N}^{T}\bm{\mathcal{G}}_{N}{\bm{u}}+\bm{h}_{N},$
where $\bm{\mathcal{G}}_{N}=\bm{P}_{N}\bm{g}_{N}\bm{\rho}$ and $\bm{h}_{N}={\bm{\varepsilon}}_{N}^{T}\bm{P}_{N}\big{(}\bm{f}_{N}+{\bm{d}_{N}}-\sum_{k=1}^{N-1}\frac{\partial\bm{a}_{N-1}}{\partial\bm{x}_{k}}\left(\bm{f}_{k}+{\bm{d}_{k}}+\bm{g}_{k}\bm{x}_{k+1}\right)-\bm{\omega}_{N-1}+\bm{g}_{N}\bm{\upsilon}\big{)}+{\bm{\varepsilon}}_{N-1}^{T}\bm{P}_{N-1}\bm{g}_{N-1}{\bm{\varepsilon}}_{N}+\frac{1}{2}{\bm{\varepsilon}}_{N}^{T}\dot{\bm{P}}_{N}{\bm{\varepsilon}}_{N}$. Similar to Step i ($i=1,\ldots,N-1$), the uncertain function $\bm{h}_{N}$ can be upper bounded by
$\bm{h}_{N}\leq\|\bm{\varepsilon}_{N}\|^{2}\theta_{N}\Phi_{N}+\frac{3}{2}$ (44)
with $\theta_{N}=\max\{{a}_{N1}^{2}a_{fN}^{2},{a}_{N1}^{2}a_{f(N-1)}^{2},{a}_{N1}^{2}\bar{g}_{N-1}^{2},{a}_{N1}^{2}\bar{g}_{N}^{2}\bar{\upsilon}^{2},\\ {a}_{N1}^{2},a_{(N-1)1}^{2}\bar{g}_{N-1}^{2},{a}_{N2}\}$ being an unknown positive constant and $\Phi_{N}={\phi}_{N1}^{2}\phi_{fN}^{2}+{\phi}_{N1}^{2}\phi_{f(N-1)}^{2}\|\frac{\partial\bm{a}_{N-1}}{\partial\bm{x}_{N-1}}\|^{2}+{\phi}_{N1}^{2}\|\frac{\partial\bm{a}_{N-1}}{\partial\bm{x}_{N-1}}\|^{2}+{\phi}_{N1}^{2}\|\bm{\omega}_{N-1}\|^{2}+{\phi}_{N1}^{2}+\phi_{(N-1)1}^{2}\|{\bm{\varepsilon}}_{N-1}\|^{2}+\frac{1}{2}{\phi}_{N2}$ denoting a known and computable scalar function.
By inserting (44) into (43), we can get
$\dot{V}_{N1}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\sum_{k=2}^{N-1}\kappa_{k}\lambda_{k}\|{\bm{\varepsilon}}_{k}\|^{2}-\sum_{k=1}^{N-1}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}$ (45)
$+\Delta_{N-1}+{\bm{\varepsilon}}_{N}^{T}\bm{\mathcal{G}}_{N}{\bm{u}}+\|\bm{\varepsilon}_{N}\|^{2}\theta_{N}\Phi_{N}+\frac{3}{2}.$
The actual control law $u$ is designed as:
$\bm{u}$ $=-\frac{\bm{A}^{T}}{\|\bm{A}\|}\left(\kappa_{N}{\bm{\varepsilon}}_{N}+{\hat{\theta}_{N}{\Phi_{N}}{{\bm{\varepsilon}}_{N}}}\right),$ (46)
$\dot{\hat{\theta}}_{N}$ $={\sigma_{N}\|{\bm{\varepsilon}}_{N}\|^{2}}\Phi_{N}-\mu_{N}\hat{\theta}_{N},\ \hat{\theta}_{N}(0)\geq 0,$ (47)
where $\kappa_{N}>0$, $\mu_{N}>0$ and $\sigma_{N}>0$ are design parameters, $\hat{\theta}_{N}$ is the estimate of $\theta_{N}$.
Remark 4
The controllability condition constructed in this paper has several merits in solving the unknown yet time-varying control gain matrix $\bm{g}_{i}$: 1) in contrast to [8, 26], no approximation methods are used to estimate $\bm{g}_{i}$, thus avoiding the singularity problem [27] that may arise when calculating the inverse of the estimated matrix; 2) it does not impose assumptions on $\bm{g}_{i}$ directly or decompose the matrix $\bm{L}$ offline to satisfy the controllability requirement, but rather, it satisfies structural assumption through the introduction of auxiliary matrix $\bm{P}_{i}$, which thus does not require any information about $\bm{g}_{i}$, and hence, can be applied to a wider variety of practical systems; and 3) avoiding using $\bm{\rho}$ to construct the Liapunov function when dealing with multiplicative actuator faults as used in [12, 15], thereby allowing for handling the more generalized intermittent faults where $\bm{\rho}$ is not satisfied to be continuous and derivable w.r.t. time.
Remark 5
As shown in (21), (32), (38) and (46), the auxiliary matrix $\bm{P}_{i}$ is not used directly in the controller, here we just use $\bm{P}_{i}$ to construct the Lyapunov function $V_{i}$. It is worth emphasizing that by embedding some “core function”-based estimators in the backstepping design procedures, the system stability obstacles posed by the unknown auxiliary matrix $\bm{P}_{i}$ and nonlinear uncertainties are elegantly handled. Specifically, the constructed controllability conditions are satisfied online by adaptive matching $\bm{P}_{i}$, where neither an extra feasibility condition on $\bm{P}_{i}$ nor tedious offline solving of the matrix $\bm{L}$ is required, thus making the proposed controllability conditions more general than those in [6, 9, 13, 10, 20, 16].
IV Stability Analysis
Now we give the stability analysis of the control laws designed in Section III-D for system (1).
Theorem 1
Consider the uncertain non-square MIMO nonlinear strict-feedback system (1) with actuator faults. Under Assumptions 1-4, if the control law (46) with the virtual controllers (21), (32) and (38) is applied, then the control objectives mentioned in Section II can be achieved.
Proof: First, we define $\tilde{\theta}_{N}=\theta_{N}-\lambda_{N}\hat{\theta}_{N}$, and then, blend such error into the complete Lyapunov function candidate such that
$V_{N}=V_{N1}+\frac{1}{2\lambda_{N}\sigma_{N}}{\tilde{\theta}_{N}}^{2},$ (48)
Differentiating (48) and using (46) yields
$\dot{V}_{N}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\sum_{k=2}^{N-1}\kappa_{k}\lambda_{k}\|{\bm{\varepsilon}}_{k}\|^{2}-\sum_{k=1}^{N-1}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}$ (49)
$-\frac{{\bm{\varepsilon}}_{N}^{T}\bm{\mathcal{G}}_{N}\bm{A}^{T}\left(\kappa_{N}{\bm{\varepsilon}}_{N}+{\hat{\theta}_{N}{\Phi_{N}}{{\bm{\varepsilon}}_{N}}}\right)}{\|\bm{A}\|}$
$+\|\bm{\varepsilon}_{N}\|^{2}\theta_{N}\Phi_{N}-\frac{1}{\sigma_{N}}{\tilde{\theta}_{N}}{\dot{\hat{\theta}}_{N}}+\Delta_{N-1}+\frac{3}{2}.$
Note that
$\bm{\mathcal{G}}_{N}\bm{A}^{T}=\,\frac{\bm{G}_{N}}{2}+\frac{\left(\bm{\mathcal{G}}_{N}\bm{A}^{T}-\bm{A}\bm{\mathcal{G}}_{N}^{T}\right)}{2},$ (50)
where $\bm{G}_{N}$ is a symmetric matrix and $\bm{\mathcal{G}}_{N}\bm{A}^{T}-\bm{A}\bm{\mathcal{G}}_{N}^{T}$ is a skew-symmetric matrix, then by Assumption 3, it holds that
${\bm{\varepsilon}}_{N}^{T}\left(\bm{\mathcal{G}}_{N}\bm{A}^{T}-\bm{A}\bm{\mathcal{G}}_{N}^{T}\right){\bm{\varepsilon}}_{N}=0,$ (51)
${\bm{\varepsilon}}_{N}^{T}\bm{G}_{N}{\bm{\varepsilon}}_{N}\geq\lambda_{N}\|\bm{A}\|\|{\bm{\varepsilon}}_{N}\|^{2}.$
Then, substituting (LABEL:sgsn2) and adaptation law (47) into (49) yields
$\dot{V}_{N}\leq$ $-\kappa_{1}\lambda_{1}\|\bm{W}{\bm{\varepsilon}}_{1}\|^{2}-\sum_{k=2}^{N}\kappa_{k}\lambda_{k}\|{\bm{\varepsilon}}_{k}\|^{2}-\sum_{k=1}^{N}\frac{\mu_{k}}{2\sigma_{k}}\tilde{\theta}_{k}^{2}$ (52)
$+\Delta_{N-1}+\frac{\mu_{N}}{2\sigma_{N}}{\theta}_{N}^{2}+\frac{3}{2}$
$\leq$ $-\Upsilon V_{N}+C,$
where $\Upsilon=\min\{\frac{2\lambda_{\textrm{min}}(\bm{W}^{T}\bm{W})\kappa_{1}\lambda_{1}}{\lambda_{\textrm{max}}(\bm{P}_{1})},\frac{2\kappa_{k}\lambda_{k}}{\lambda_{\textrm{max}}(\bm{P}_{N})},\mu_{k}\}>0$ and $C=\Delta_{N-1}+\frac{\mu_{N}}{2\sigma_{N}}{\theta}_{N}^{2}+\frac{3}{2}>0$.
Now we first prove that all signals in the closed-loop systems are bounded. It is seen from (52) that $V_{N}(t)\leq V_{N}(0)+\frac{C}{\Upsilon}=\Xi$, it follows that $V_{N}(t)$ is bounded for $t\in[0,\tau_{\textrm{max}})$. According to the definition of $V_{N}(t)$, it follows that ${\bm{\varepsilon}}_{i}$ and $\tilde{\theta}_{i}$ are bounded and satisfy ${\bm{\varepsilon}}_{1}\leq\sqrt{\frac{2\lambda_{\textrm{min}}(\bm{W}^{T}\bm{W})\Xi}{\lambda_{\textrm{max}}(\bm{P}_{1})}}$, ${\bm{\varepsilon}}_{k}\leq\sqrt{\frac{2\Xi}{\lambda_{\textrm{max}}(\bm{P}_{k})}}$ ($k=2,\ldots,N$), and $\tilde{\theta}_{i}\leq\sqrt{2\sigma_{i}\Xi}$ for $t\in[0,\tau_{\textrm{max}})$ with ${\lambda_{\textrm{max}}(\bm{P}_{k})}>0$, which further indicates that the parameter estimate $\hat{\theta}_{i}$ is bounded over the interval $t\in[0,\tau_{\textrm{max}})$. Since ${\bm{\varepsilon}}_{1}$ is bounded, and using the properties of $s_{j}$, it is shown that $-1\leq-\underline{\delta}_{j}<-\underline{\delta}_{1j}\leq\zeta_{j}(t)\leq\bar{\delta}_{1j}<\bar{\delta}_{j}\leq 1$ for $t\in[0,\tau_{\textrm{max}})$. As $\zeta_{j}=\frac{\eta_{j}}{\varphi_{j}}$ and $0<\varphi_{jf}<\varphi_{j}(t)\leq\varphi_{j0}\leq 1$, it is proved that there exist some constants $\underline{\eta}_{j}$ and $\bar{\eta}_{j}$ so that $-1<\underline{\eta}_{j}\leq\eta_{j}\leq\bar{\eta}_{j}<1$, which further implies from $e_{j}=\frac{l_{j}\eta_{j}}{\sqrt{1-\eta_{j}^{2}}}$ that the tracking error $e_{j}$ is bounded for $t\in[0,\tau_{\textrm{max}})$. As $e_{j}=x_{1j}-y_{dj}$, one has $x_{1j}$ is bounded since $y_{dj}$ is bounded for $t\in[0,\tau_{\textrm{max}})$. Hence, it is trivial to prove that $0<\underline{\mu}_{j}\leq\mu_{j}\leq\bar{\mu}_{j}$ and $0<\underline{r}_{j}\leq r_{j}\leq\bar{r}_{j}$, $0<\underline{w}_{j}\leq w_{j}\leq\bar{w}_{j}$ for $t\in[0,\tau_{\textrm{max}})$, with $\underline{\mu}_{j}$, $\bar{\mu}_{j}$, $\underline{r}_{j}$, $\bar{r}_{j}$,$\underline{w}_{j}$, $\bar{w}_{j}$ being some positive constants, which indicates that $\bm{W}$ and $\bm{V}$ are bounded. Since $x_{1}$ is bounded, it is shown from Assumption 1 and 4 that $\Phi_{1}$ is bounded, and thus $\bm{a}_{1}$ and $\dot{\hat{\theta}}_{1}$ are bounded by (21) and (22). Following this line of argument, $\bm{a}_{i}$ ($i=2,\ldots,N-1$), $\bm{u}$, $\dot{\hat{\theta}}_{k}$ ($k=2,\ldots,N$) and $\bm{x}_{k}$ are bounded over the interval $t\in[0,\tau_{\textrm{max}})$. Note that the boundedness of all signals relies on the design parameters, yet is independent of $\tau_{\textrm{max}}$, which reveals that $\tau_{\textrm{max}}$ can be extended to $\tau_{\textrm{max}}=\infty$. Thus, all signals in the closed-loop systems are bounded on $t\in[0,\infty)$.
Next we prove that the tracking error is always within the predefined region for $t\in[0,\infty)$. As $s_{j}$ is bounded for $t\in[0,\infty)$, according to Lemma 1, it is ensured that $\mathcal{H}(-\underline{\delta}_{j}\varphi_{j}(t))<e_{j}(t)<\mathcal{H}(\bar{\delta}_{j}\varphi_{j}(t))$, $j=1,2,\ldots,m$. The proof is completed. $\square$
Remark 6
It can be seen that the system (1) with output constraints is transformed to an unconstrained system (14) by introducing a unified prescribed performance function $\mathcal{H}$. Then, by applying the proposed virtual controllers (21), (32), (38) and actual controller (46) to ensure the stability of the transformed system, the predefined performance constraint on the tracking error is indirectly implemented. It is worth noting that by adjusting the key design parameters, the global performance and semi-global yet asymmetric performance can be achieved without changing the control structure. Furthermore, the unexpected actuator faults can be compensated automatically without resorting to any additional fault detection and diagnosis module.
Remark 7
The controllability condition and controller constructed in this paper can be flexibly applied to a class of square and non-square systems. Specifically, square and non-square systems can be handled by simply choosing $\bm{A}=\bm{I}_{n}$ and $\bm{A}=[\bm{I}_{n},\bm{\Lambda}_{1},\ldots,\bm{\Lambda}_{(m-n)}]$, respectively. Moreover, since the auxiliary matrix $\bm{P}_{i}$ is adaptively compensated and does not use in the controller, even if $\bm{g}_{i}$ satisfies the original strong controllability condition in [9], it is sufficient to compensate $\bm{P}_{i}$ as a identity matrix $\bm{I}_{n}$ without any additional operation.
V Illustrative Example
In order to validate the effectiveness, we implement the proposed control approach into a spacecraft actuated by four RWs [19], for which the dynamics can be derived as
$\dot{\bm{x}}=\bm{f}(\bm{x})+\bm{g}(\bm{x},t)\bm{u}_{a}+\bm{d}(t),$ (53)
with $\bm{x}=\bm{\omega}$, $\bm{f}(\bm{x})=-\bm{J}^{-1}\bm{\omega}^{\times}\bm{J}\bm{\omega}$, $\bm{g}(\bm{x},t)=\bm{J}^{-1}\bm{D}$, $\bm{u}_{a}=\bm{\tau}$ and $\bm{d}(t)=\bm{J}^{-1}\bm{d}_{0}(t)$, where $\bm{\omega}\in\mathbb{R}^{3}$ is reference angular velocity, $\bm{J}\in\mathbb{R}^{3\times 3}$ is the total inertia matrix, $\bm{D}\in\mathbb{R}^{3\times 4}$ (4 is the number of RWs) is the RWs configuration matrix with full-row rank, while $\bm{\tau}\in\mathbb{R}^{4}$ and $\bm{d}_{0}(t)\in\mathbb{R}^{3}$ denote the control and bounded disturbance torques, respectively. Obviously, this is an non-square system with four control inputs and three outputs.
Figure 1: The evolution of $e_{j}$ with different initial conditions, $i=1,2,3$.
Figure 2: Control input signals under $\bm{x}_{1}(0)=[1,0.5,0.5]^{T}$.
Figure 3: The evolution of $e_{j}$ under $\underline{\delta}_{j}=0.8,\ \bar{\delta}_{j}=1$, $i=1,2,3$.
Figure 4: The evolution of $e_{j}$ under $\underline{\delta}_{j}=1,\ \bar{\delta}_{j}=0.7$, $i=1,2,3$.
The unknown inertia matrix $\bm{J}$ is given by $\bm{J}=\bm{J}_{0}+\bm{J}_{u}$, where $\bm{J}_{0}$ is the nominal part set as
$\bm{J}_{0}=\begin{bmatrix}40&1.2&0.9\\ 1.2&17&1.4\\ 0.9&1.4&15\end{bmatrix}\textrm{kg}\cdot\textrm{m}^{2},$
while $\bm{J}_{u}$ is the uncertain part given by $\bm{J}_{u}=\textrm{diag}\{0.2e^{-0.2t},\\ 2e^{-0.1t},3e^{-0.1t+1}\}\textrm{kg}\cdot\textrm{m}^{2}$. The configuration matrix is set as
$\bm{D}=\begin{bmatrix}1&0&0&{1}/{\sqrt{3}}\\ 0&1&0&{1}/{\sqrt{3}}\\ 0&0&1&{1}/{\sqrt{3}}\end{bmatrix}.$
The PLOE fault profiles are outlined as follows:
$\bm{\rho}(t)=\begin{cases}\textmd{diag}\big{\{}0.9+0.1\sin(t),1-0.2\tanh(t),\\ \ 0.9+0.1\cos(t),0.8-0.2\sin(t)\big{\}},\ \ \,t\in(0,5]\\ \textmd{diag}\big{\{}0.8+0.2\sin(t),0.8-0.2\tanh(t),\\ \ 0.8+0.2\cos(t),0.8-0.1\tanh(t)\big{\}},\ t\in(5,\infty)\end{cases}$
$\bm{\upsilon}(t)=[0.01\tanh(2t),0.01\cos(t),0.01\sin(3t),0.01\cos(2t)]^{T}.$
The key design parameters are selected as $\underline{\delta}_{j}=\bar{\delta}_{j}=1$, $j=1,2,3$. The control parameters are given as: $\kappa_{1}=1$, $\sigma_{1}=0.01$ and $\mu_{1}=0.1$. The matrix $\bm{A}$ is setting as
$\bm{A}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&1\end{bmatrix}.$
The rate function is chosen as $\beta(t)=\exp(-0.9t)\cos^{2}(t)$, and $\varphi_{j0}=1$, $\varphi_{jf}=0.1$ and $l=0.9936$, $j=1,2,3$. The initial conditions are set as: $\bm{x}_{1}(0)=[-0.5,-0.5,-0.5]^{T}$(rad) and $[1,0.5,0.5]^{T}$(rad), $\hat{\theta}_{1}=0$. The simulation results are given in Figs.1 and 2. Fig.1 shows that the output tracking errors $e_{j}$, $j=1,2,3$, evolve within the prescribed performance bounds ($\mathcal{H}(-\varphi_{j}(t)),\mathcal{H}(\varphi_{j}(t))$). Fig.2 shows the boundedness of the control signal $u$.
To verify that the proposed control is able to achieve the asymmetric performance behaviors, we select the key design parameters as $\underline{\delta}_{j}=0.8,\ \bar{\delta}_{j}=1$ and $\underline{\delta}_{j}=1,\ \bar{\delta}_{j}=0.7$, $j=1,2,3$. The initial conditions are set as: $\bm{x}_{1}(0)=[-0.5,-0.5,-0.5]^{T}$(rad) and $[1,0.5,0.5]^{T}$(rad). Other parameters and conditions are kept consistent with the previous simulations. The responses of tracking errors are depicted in Fig.3 and Fig.4, which shows that the tracking errors are always evolve within the prescribed performance bounds, in line with the theoretical analysis.
VI Conclusion
In this paper, a unified prescribed performance tracking control method with generalized controllability condition is proposed for a class of non-square nonlinear strict-feedback systems with actuator faults. By using a practical matrix decomposition and resorting to some suitable auxiliary matrices, a more relaxed controllability condition for non-square systems with actuator faults is meticulously constructed. The proposed method can achieve global performance and semi-global performance yet asymmetric performance behaviors by choosing key design parameters without changing the control structure. By constructing some augmented Lyapunov functions using the auxiliary matrices instead of the actuation effectiveness matrix, not only the controllability of the closed-loop system is guaranteed, but also the stringent constraints on the original gain matrix and actuation effectiveness matrix are eliminated. In the future, we aim to generalize the method to non-square systems in the under-actuated case.